# Labeled Data Selection for Category Discovery

### v1

@@ Abstract
Category discovery methods aim to find novel categories in unlabeled visual data. At training time, a set of labeled and unlabeled images are provided, where the labels correspond to the categories present in the images. The labeled data provides guidance during training by indicating what types of visual properties and features are relevant for performing discovery in the unlabeled data. As a result, changing the categories present in the labeled set can have a large impact on what is ultimately discovered in the unlabeled set. Despite its importance, the impact of labeled data selection has not been explored in the category discovery literature to date. We show that changing the labeled data can significantly impact discovery performance. Motivated by this, we propose two new approaches for automatically selecting the most suitable labeled data based on the similarity between the labeled and unlabeled data. Our observation is that, unlike in conventional supervised transfer learning, the best labeled is neither too similar, nor too dissimilar, to the unlabeled categories. Our resulting approaches obtains state-of-the-art discovery performance across a range of challenging fine-grained benchmark datasets.
@@ 1 Introduction
On first appearance, the availability of labeled data may appear to only be a minor difference from the unsupervised clustering setting. However, the labeled data plays an important role for discovery algorithms. It provides guidance during training regarding the types of visual features that are deemed important and relevant for defining categories in the unlabeled data. For example, suppose one was provided with a dataset that contains images captured in nature. If the labeled data only contained images of different species of birds, this strongly limits the discovery method to only be able to find ‘bird-like’ concepts in the unlabeled data. Similarly, if the labeled data instead only contained images of plant species, it would restrict the method to discover ‘plant-like’ concepts.
@@
As a result, the role and selection of labeled data is an important question in category discovery. However, until now, this problem has been overlooked in the discovery literature, which has instead focused on developing new methods that can be applied to off-the-shelf datasets with already provided labeled and unlabeled splits. In the context of supervised classification, [9] showed that the choice of pre-training source data can have a large impact on downstream fine-tuned target performance. Motivated by this, they proposed a new approach for selecting a source training set that is most similar to a target domain of interest by measuring the similarity between the source and target domains. Their observation is that the ideal source dataset for supervised classification is the one that is most similar to the target domain. In this work, we show that this is not the case for category discovery.
@@
In discovery, we demonstrate that the similarity between the labeled categories and the unlabeled data exposes a surprising, and at first perhaps counter-intuitive, relationship. Intuitively, if the categories in the labeled data are too dissimilar to the ones in the unlabeled data, discovery fails, i.e. labeled images of insects do not provide useful cues for discovering species of birds. However, if the labeled categories are too similar to the unlabeled ones this also dramatically hurts discovery performance, i.e. if the labeled and unlabeled datasets both contain very visually similar, but distinct, fine-grained bird categories, discovery methods are unable to separate the two sets of similar categories. As a result, there appears to be a ‘sweet spot’ in the middle of both extremes, whereby the labeled images should be related enough to provide useful context, but not too similar as to be overly difficult to separate. Exploiting this observation to select a subset of labeled categories for training category discovery methods results in significantly enhanced discovery performance (see Fig. 1).
@@
We introduce a new labeled dataset selection framing of the generalized category discovery problem. At training time we are presented with an unlabeled target dataset in which we want to discover novel categories. In addition, we are given a labeled source dataset where the goal is to automatically select a subset from it to train the discovery algorithm. Importantly, this labeled dataset can potentially contain images from diverse categories, many of which may not be relevant for the unlabeled target. We propose two new unsupervised selection methods that assign weight to the labeled data during the training process so that labeled data that is too similar or too dissimilar to the target data is down-weighted and thus has less influence on the learned representation.
@@ 2.1 Category discovery
In novel category discovery (NCD), the unlabeled data is assumed to not contain any images from the categories in the labeled set [20, 18, 19, 53, 15, 56]. This is a overly simplifying assumption, and in the more realistic generalized category discovery (GCD) setting the unlabeled data can contain images from both previously seen and novel categories [45, 5, 49, 33, 14, 51, 47]. Both of these discovery paradigms operate under a closed set assumption, i.e. there are no novel categories in the test set that are not in the labeled or unlabeled sets. On-the-fly category discovery (OCD) attempts to handle this specific challenge [12, 34]. Finally, there is a recent set of methods that explore discovery in the online learning setting [52, 24, 54]. Here, the goal is to not only discover new concepts, but also not to forget previously seen ones.
@@ 3.3 Data selection
Our central hypothesis is that not all labeled data is equally useful for performing category discovery. To address this, we aim to ‘select’ a subset of the entire labeled source dataset to use for training the discovery model (e.g. SimGCD). We perform data selection by generating instance specific weights for each data point in the labeled source dataset. Specifically, the selection process aims to generate a weight $\omega_{i}\in[0,+\infty)$ for each data point $\mathbf{x}_{i}$ in the labeled dataset which is used to weight some components of the training loss in Eqs. 4 and 2. Thus, data points with weight 00 are effectively discarded during training, and non-zero weighted data points influence training based on the magnitude of their weight. Similarly, using a weight of $1$ for all data points is equivalent to running off-the-shelf discovery with no weighting. Next we review a related supervised selection approach proposed for improving transfer learning and then introduce two new unsupervised selection methods.
@@
Figure 2: (Left) The beta distribution used to generate weights for the labeled data. (Right) Overview of our labeled data selection process for category discovery. The labeled data is weighted based on the distance to the unlabeled data. The weight is used to change the influence of labeled categories during training. The green and orange arrows on the left panel denote the distance between labeled to unlabeled data for two instances illustrated using same color in the first step on the right panel.
@@ 4.1 Implementation details
Evaluation protocol. To evaluate model performance, we adopt the commonly used clustering accuracy metric [45]. During evaluation, given the ground truth labels $y$ and the model predicted labels $\tilde{y}$, we calculate the accuracy as $\text{ACC}=\frac{1}{M}\sum_{i=1}^{M}\mathds{1}(y^{*}_{i}=\mathcal{P}(\tilde{y}_{i}))$, where $M=|\mathcal{D}^{u}|$, and $\mathcal{P}$ is the optimal permutation that matches the predicted cluster assignments to the ground truth class labels. Additionally, we report ‘All’, ‘New’, and ‘Old’ performance as is common in the literature. ‘All’ is for the performance on all clusters, and ‘New’ and ‘Old’ denote the performance on novel and seen categories respectively. We report the average of three different runs in each table.
@@ 4.2 Evaluation
Perhaps surprisingly, the source split that performs the best (‘Medium’) is a subset of the source data that contains categories that are not too similar nor too dissimilar. Using the ‘Medium’ labeled data alone can match, or even outperform, using all labeled data in some cases (i.e. on CUB and Aircraft), despite using only one-third of the data. By taking a closer look at the source of errors in Fig. 3, we identify that when the model is using the ‘Similar’ splits, the majority of errors stem from misclassifying the novel categories to previously seen ones. When the model is trained with ‘Dissimilar’ splits, the error is mainly due to misclassifying the novel categories to incorrect novel clusters. When the labeled split comes from Out-of-Domain (OOD) insects, the error is almost entirely from misclassifying novel categories as incorrect novel clusters. These results indicate that when the model is trained with ‘Dissimilar’ or OOD data, it can learn to predict the novel categories as new clusters, but cannot learn a discriminative enough representation to be able to separate the novel categories apart. On the other hand, when the model is trained with ‘Similar’ data, it struggles to separate novel categories from seen ones, resulting in a higher error when misclassifying novel categories as seen.
Figure 3: Percentage of errors made on CUB when using ‘Similar’, ‘Medium’, ‘Dissimilar’, and ‘OOD’ labeled training subsets. ‘Misclassified as New’ means that an example from a ‘New’ category is assigned to another ‘New’ one, and ‘Misclassified as Old’ means it is assigned to an ‘Old’ category. Absolute scores are in Tab. 1.
Table 2: Category discovery results on the Semantic Shift Benchmark [46]. Here, labeled data is selected from all the datasets combined, i.e. we combined all the labeled data from the four datasets and the OOD insects.
@@
Table 3: Discovery results on NABirds [42], Herbarium-19 [40], and ImageNet [35]. The source data consists of these three datasets and the source data used in Tab. 4.
@@ 4.3 Ablations
Different of $\alpha,\beta$ values. In Sec. 3.3, we showed that by changing the value of $\alpha$ and $\beta$, we can control the shape of the curve for generating data weighs thus interpolating between different strategies. In Tab. A2 we observe that when setting the values to $\alpha=5$ and $\beta=1$, to favor more similar samples like in [9], the performance drops. Setting both $\alpha$ and $\beta$ values to larger than $1$ yields consistent performance improvements, and our default choice of $\alpha=\beta=5$ yields consistently strong performance across the three datasets.
Hard threshold cutting. In Sec. 3.3, we discussed using the weight generated by Eq. 13 to weigh the influence of each example. Tab. A2 shows the ablation of thresholding the generated weight to $\{0,1\}$. By thresholding the weight $\omega_{i}$ of an instance $\mathbf{x}_{i}$ to 00, we can effectively remove that sample from the training dataset. Thus the training time can be reduced. These results demonstrate that while using the soft weight performs the best, removing some examples from training not only does not hurt performance but improves it to some extent.
@@ 4.4 Limitations
One limitation of our proposed data selection method is that it requires specifying some hyperparameters (e.g. the parameters of the beta distribution). Notably, we use the same hyperparameters across all datasets and demonstrate the robustness of our method to different choices. There is also an additional cost incurred by our approach associated with computing the training weights. However, this is negligible compared to the overall cost of training. Like all discovery methods, if the resulting discovered categories were not verified by a human expert, there could be potential negative impact of our work if it was applied in safety-critical applications.
@@ 5 Conclusion
Acknowledgements. This work was in part supported by a Royal Society Research Grant and the Edinburgh–Copenhagen Strategic Partnership Seed-Fund.

### v2

@@ Abstract
Visual category discovery methods aim to find novel categories in unlabeled visual data. At training time, a set of labeled and unlabeled images are provided, where the labels correspond to the categories present in the images. The labeled data provides guidance during training by indicating what types of visual properties and features are relevant for performing discovery in the unlabeled data. As a result, changing the categories present in the labeled set can have a large impact on what is ultimately discovered in the unlabeled set. Despite its importance, the impact of labeled data selection has not been explored in the category discovery literature to date. We show that changing the labeled data does indeed significantly impact discovery performance. Motivated by this, we propose two new approaches for automatically selecting the most suitable labeled data based on the similarity between the labeled and unlabeled data. Our observation is that, unlike in conventional supervised transfer learning, the most informative labeled data is neither too similar nor too dissimilar, to the unlabeled categories. Our resulting approaches obtain state-of-the-art discovery performance across a range of methods and challenging fine-grained benchmark datasets.
@@ 1 Introduction
On first appearance, the availability of labeled data may appear to only be a minor difference from traditional unsupervised clustering. However, the labeled data plays an important role for discovery algorithms. It provides guidance during training regarding the types of visual features that are deemed important and relevant for defining categories in the unlabeled data. For example, suppose one was provided with a dataset that contains images captured in nature. If the labeled data only contained images of different species of birds, this strongly limits the discovery method to only be able to find ‘bird-like’ concepts in the unlabeled data. Similarly, if the labeled data instead only contained images of plant species, it would restrict the method to discover ‘plant-like’ concepts.
@@
As a result, the role and selection of labeled data is an important question in category discovery. However, until now, this problem has been overlooked in the discovery literature, which has instead focused on developing new methods that can be applied to off-the-shelf datasets with already provided labeled and unlabeled splits. In the context of supervised classification, [9] showed that the choice of labeled pre-training source data can have a large impact on downstream fine-tuned target performance. Motivated by this, they proposed an approach for selecting a source training set that is most similar to a target domain of interest by measuring the similarity between the source and target domains. Their observation is that the ideal source dataset for supervised classification is the one that is most similar to the target domain. In this work, we show that this is not the case for category discovery.
@@
In discovery, we demonstrate that the similarity between the labeled categories and the unlabeled data exposes a surprising, and at first perhaps counter-intuitive, relationship. Intuitively, if the categories in the labeled data are too dissimilar to the ones in the unlabeled data, discovery fails, i.e. labeled images of insects do not provide useful cues for discovering species of birds. However, if the labeled categories are too similar to the unlabeled ones this also dramatically hurts discovery performance, i.e. if the labeled and unlabeled datasets both contain very visually similar, but distinct, fine-grained bird categories, discovery methods are unable to separate the two sets of similar categories. As a result, there appears to be a ‘sweet spot’ in the middle of both extremes, whereby the labeled images should be related enough to provide useful context, but not too similar as to be overly difficult to separate. We show that exploiting this observation to select a subset of labeled categories for training category discovery methods results in significantly enhanced discovery performance (see Fig. 1).
@@
We introduce a new labeled dataset selection framing of the generalized category discovery problem. At training time we are presented with an unlabeled target dataset in which we want to discover novel categories. In addition, we are given a labeled source dataset where the goal is to automatically select a subset from it to train the discovery algorithm. Importantly, this labeled dataset can potentially contain images from diverse categories, many of which may not be relevant to the unlabeled target. We propose two new unsupervised selection methods that assign weight to the labeled data during the training process so that labeled data that is too similar or too dissimilar to the target data is down-weighted and thus has less influence on the learned representation.
@@ 2.1 Category discovery
In novel category discovery (NCD), the unlabeled data is assumed to not contain any images from the categories in the labeled set [20, 18, 19, 53, 15, 56]. This is a overly simplifying assumption, and in the more realistic generalized category discovery (GCD) setting the unlabeled data can contain images from both previously seen and novel categories [45, 5, 49, 33, 14, 51, 47]. Both of these discovery paradigms operate under a closed set assumption, i.e. there are no novel categories in the test set that are not in the labeled or unlabeled sets. On-the-fly category discovery (OCD) attempts to handle this specific challenge [12, 34]. Finally, there is a recent set of methods that explore discovery in the online learning setting [52, 24, 54]. Here, the goal is to not only discover new concepts but also not to forget previously seen ones.
@@ 3.3 Data selection
Our central hypothesis is that not all labeled data is equally useful for performing category discovery. To address this, we aim to ‘select’ a subset of the entire labeled source dataset to use for training the discovery model (e.g. SimGCD). We perform data selection by generating instance-specific weights for each data point in the labeled source dataset. Specifically, the selection process aims to generate a weight $\omega_{i}\in[0,+\infty)$ for each data point $\mathbf{x}_{i}$ in the labeled dataset which is used to weight some components of the training loss in Eqs. 4 and 2. Thus, data points with weight 00 are effectively discarded during training, and non-zero weighted data points influence training based on the magnitude of their weight. Similarly, using a weight of $1$ for all data points is equivalent to running off-the-shelf discovery with no weighting. Next, we review a related supervised selection approach proposed for improving transfer learning and then introduce two new unsupervised selection methods.
@@
Figure 2: (Left) The beta distribution is used to generate weights for the labeled data. (Right) Overview of our labeled data selection process for category discovery. The labeled data is weighted based on the distance to the unlabeled data. The weight is used to change the influence of labeled categories during training. The green and orange arrows on the left panel denote the distance between labeled to unlabeled data for two instances illustrated using same color in the first step on the right panel.
@@ 4.1 Implementation details
Evaluation protocol. To evaluate model performance, we adopt the commonly used clustering accuracy metric [45]. During the evaluation, given the ground truth labels $y$ and the model predicted labels $\tilde{y}$, we calculate the accuracy as $\text{ACC}=\frac{1}{M}\sum_{i=1}^{M}\mathds{1}(y^{*}_{i}=\mathcal{P}(\tilde{y}_{i}))$, where $M=|\mathcal{D}^{u}|$, and $\mathcal{P}$ is the optimal permutation that matches the predicted cluster assignments to the ground truth class labels. Additionally, we report ‘All’, ‘New’, and ‘Old’ performance as is common in the literature. ‘All’ is for the performance on all clusters, and ‘New’ and ‘Old’ denote the performance on novel and seen categories respectively. We report the average of three different runs in each table.
@@ 4.2 Evaluation
Figure 3: Percentage of errors on three datasets when using ‘Similar’, ‘Medium’, ‘Dissimilar’, and ‘OOD’ labeled training sets. ‘Misclassified as New’ means that an example from a ‘New’ category is assigned to another ‘New’ one, and ‘Misclassified as Old’ means it is assigned to an ‘Old’ one.
Perhaps surprisingly, the source split that performs the best (‘Medium’) is a subset of the source data that contains categories that are not too similar nor too dissimilar. Using the ‘Medium’ labeled data alone can match, or even outperform, using all labeled data in some cases (i.e. on CUB and Aircraft), despite using only one-third of the data. By taking a closer look at the source of errors in Fig. 3, we identify that when the model is using the ‘Similar’ splits, the majority of errors stem from misclassifying the novel categories to previously seen ones. When the model is trained with ‘Dissimilar’ splits, the error is mainly due to misclassifying the novel categories to incorrect novel clusters. When the labeled split comes from Out-of-Domain (OOD) insects, the error is almost entirely from misclassifying novel categories as incorrect novel clusters. These results indicate that when the model is trained with ‘Dissimilar’ or OOD data, it can learn to predict the novel categories as new clusters, but cannot learn a discriminative enough representation to separate the novel categories apart. On the other hand, when the model is trained with ‘Similar’ data, it struggles to separate novel categories from seen ones, resulting in a higher error when misclassifying novel categories as seen.
Table 2: Category discovery results on the Semantic Shift Benchmark [46]. Here, labeled data is selected from all the datasets combined, i.e. we combined all the labeled data from the four datasets and the OOD insects. Performance change with respect to no-selection in presented in parentheses.
@@
Table 3: Discovery results on NABirds [42], Herbarium-19 [40], and ImageNet [35]. The source data consists of these three datasets plus the source data used in Tab. 4.
@@ 4.3 Ablations
Different of $\alpha,\beta$ values. In Sec. 3.3, we showed that by changing the value of $\alpha$ and $\beta$, we can control the shape of the curve for generating data weighs thus interpolating between different strategies. In Tab. 6 we observe that when setting the values to $\alpha=5$ and $\beta=1$, to favor more similar samples like in [9], the performance drops. Setting both $\alpha$ and $\beta$ values to larger than $1$ yields consistent performance improvements, and our default choice of $\alpha=\beta=5$ yields consistently strong performance across the three datasets.
Hard threshold cutting. In Sec. 3.3, we discussed using the weight generated by Eq. 13 to weigh the influence of each example. Tab. 6 shows the ablation of thresholding the generated weight to $\{0,1\}$. By thresholding the weight $\omega_{i}$ of an instance $\mathbf{x}_{i}$ to 00, we can effectively remove that sample from the training dataset. Thus the training time can be reduced. These results demonstrate that while using the soft weight performs the best, removing some examples from training not only does not hurt performance but improves it to some extent.
@@ 4.4 Limitations
One limitation of our proposed data selection method is that it requires specifying some hyperparameters (e.g. the parameters of the beta distribution). Notably, we use the same hyperparameters across all datasets and demonstrate the robustness of our method to different choices. There is also an additional cost incurred by our approach associated with computing the source data training weights. However, this is negligible compared to the overall cost of training. Like all discovery methods, if the resulting discovered categories were not verified by a human expert, there could be potential negative impact of our work if it was applied in safety-critical applications.
@@ 5 Conclusion
Acknowledgements. This work was in part supported by a Royal Society Research Grant, the Edinburgh–Copenhagen Strategic Partnership Seed-Fund, and the Pioneer Centre for AI, DNRF grant number P1.